# ProSparse: Introducing and Enhancing Intrinsic Activation Sparsity within Large Language Models

### v4

@@ 3.2 ProSparse
Concretely, for the warmup stage, we set $\lambda$ to a constant value, which is relatively small to prevent radical activation distribution shifts and introduce higher preliminary sparsity. Next, for each of the remaining stages (called incremental stages), $\lambda$ is scheduled to increase along a smooth sine curve from a trough value to a peak value. Inspired by the cosine annealing scheduler for learning rates [46], we choose the sine function for $\lambda$ increase owing to its special trend. Specifically, the sine function has small derivatives near the trough and the peak, thereby $\lambda$ will not increase radically around these two points. This provides the LLMs with more time to adapt the activation distributions to the newly increased $L_{1}$ regularization. Notably, each stage is accompanied by certain steps of training. The step numbers and peak $\lambda$ values are chosen according to the target sparsity and stability requirements.
{wrapfloat}
@@
algorithm[18]r0.5 Progressive sparsity regularization
@@
Activation Threshold Shifting
As demonstrated by recent works, there exist considerable amounts of non-zero low elements in the activation outputs, which have little influence on final results and thus can be pruned for higher sparsity [88]. Therefore, we convert the ReLU into FATReLU [38] by shifting the activation threshold, i.e.,
where $t>0$ is a positive threshold. As long as $t$ is properly chosen (see Appendix J), FATReLU can increase sparsity with negligible losses [88].
3.3 Practical Inference Acceleration
To go further beyond former theoretical acceleration analyses based on FLOPS (Floating Point Operations Per Second) [50] and establish the practical value of ProSparse, in this section, we discuss how to realize inference acceleration with sparsely activated LLMs on real GPU hardware and how to evaluate the practical acceleration effects. We consider two categories of acceleration algorithms based on activation sparsity: the approximate algorithm and the accurate algorithm.
Approximate Acceleration Algorithms
Utilizing activation sparsity, recent approximate acceleration algorithms predominantly rely on activation predictors, typically small neural networks, to forecast the activation distributions indicated by the sparse intermediate outputs $\mathbf{x}_{1}$ given a specific input $\mathbf{x}$ [44, 66]. In this way, they can make wiser hardware allocation or computation policies to avoid resource waste on weakly-contributed parameters. However, their efficiency and accuracy largely depend on the predictors’ performance, and invalid predictions can cause suboptimal hardware allocation or even inference inaccuracy. Therefore, to gain more practical acceleration effects from approximate algorithms, both high activation sparsity and predictability are indispensable.
To this end, we focus on three metrics for acceleration effect analysis: the activation recall (hereinafter abbreviated as recall), the predicted sparsity, and the inference speed. The former two metrics evaluate the performance of activation predictors as well as the activation predictability of a sparse LLM [88].

### v5

@@ 3.2 ProSparse
Concretely, for the warmup stage, we set $\lambda$ to a constant value, which is relatively small to prevent radical activation distribution shifts and introduce higher preliminary sparsity. Next, for each of the remaining stages (called incremental stages), $\lambda$ is scheduled to increase along a smooth sine curve from a trough value to a peak value. Inspired by the cosine annealing scheduler for learning rates Loshchilov and Hutter (2016), we choose the sine function owing to its special trend, as the small derivatives near its troughs and peaks can make $\lambda$ not increase radically around these two points. This provides the LLMs with more time to adapt the activation distributions to the newly increased $L_{1}$ regularization. Notably, each stage is accompanied by certain steps of training. The step numbers and peak $\lambda$ values are chosen according to the target sparsity and stability.
Activation Threshold Shifting
As demonstrated by recent works, there exist considerable amounts of non-zero low elements in the activation outputs, which have little influence on final results and thus can be pruned for higher sparsity Zhang et al. (2024). Therefore, we convert the ReLU into FATReLU Kurtz et al. (2020) by shifting the activation threshold, i.e.,
where $t>0$ is a positive threshold. As long as $t$ is properly chosen (see Appendix O), FATReLU can increase sparsity with negligible losses Zhang et al. (2024).
3.3 Practical Inference Acceleration
To go beyond theoretical analyses based on FLOPS (Floating Point Operations Per Second) Mirzadeh et al. (2023) and establish the practical value of ProSparse, we discuss how to realize inference acceleration with sparsely activated LLMs on real hardware and how to evaluate the practical acceleration effects. We consider two categories of acceleration algorithms based on activation sparsity: approximate algorithms and accurate algorithms.
Approximate Acceleration Algorithms
Utilizing activation sparsity, recent approximate acceleration algorithms predominantly rely on activation predictors, typically small neural networks, to forecast the activation distributions indicated by the sparse intermediate outputs $\mathbf{x}_{1}$ given a specific input $\mathbf{x}$ Liu et al. (2023); Song et al. (2023). In this way, they can make wiser hardware allocation or computation policies to avoid resource waste on weakly-contributed parameters. However, their efficiency and accuracy largely depend on the predictors’ performance, and invalid predictions can cause suboptimal hardware allocation or even inference inaccuracy. Therefore, to gain more practical acceleration effects from approximate algorithms, both high activation sparsity and predictability are indispensable.
To this end, we focus on three metrics for acceleration analysis: the activation recall, the predicted sparsity, and the inference speed. The former two metrics evaluate the performance of activation predictors as well as the activation predictability of a sparse LLM Zhang et al. (2024). For inference speed, we adopt PowerInfer Song et al. (2023), a state-of-the-art approximate algorithm to measure practical speedup ratios. Refer to Appendix C for more related introductions and the detailed approach to calculating these metrics.
Accurate Acceleration Algorithms
Table 1: The overall experimental results with the comparison of activation sparsity (%) and downstream performance (%). “LLaMA2” and “MiniCPM” refer to the original Swish-activated LLaMA2 Touvron et al. (2023b) and MiniCPM Hu et al. (2024) versions respectively. “ProSparse-7B”, “ProSparse-13B”, and “ProSparse-1B” denote the ProSparse versions without activation threshold shifting.
Generation Reasoning Comprehension Performance Sparsity
LLaMA2-7B 16.37 69.59 61.87 12.96 44.45 32.96 27.53 37.96 -
ReluLLaMA-7B 15.85 69.64 70.54 5.84 38.64 35.07 27.73 37.62 66.98
ProSparse-7B 19.47 66.29 63.33 12.74 45.21 33.59 27.55 38.31 88.11
ProSparse-7B 19.42 66.27 63.50 12.13 45.48 34.99 27.46 38.46 89.32
LLaMA2-13B 20.19 72.58 71.55 22.21 54.69 37.89 29.33 44.06 -
ReluLLaMA-13B 20.19 70.44 73.29 18.50 50.58 37.97 28.22 42.74 71.56
ProSparse-13B 29.03 69.75 67.54 25.40 54.78 40.20 28.76 45.07 87.97
ProSparse-13B 28.42 69.76 66.91 26.31 54.35 39.90 28.67 44.90 88.80
MiniCPM-1B 36.85 63.67 60.90 35.48 50.44 35.03 28.71 $4$
ProSparse-1B 41.38 64.55 60.69 34.72 49.36 34.04 28.27 44.72 86.25
ProSparse-1B 42.04 64.37 60.73 34.57 49.51 34.08 27.77 44.72 87.89
Concretely, we reorganize a ReLU-activated gated FFN into three major steps and our two operators are responsible for the step (2) and (3) respectively: (1) A dense matrix-vector multiplication operator $\mathbf{x}\mathbf{W}_{s}^{T}$ directly supported by vendor libraries such as cuBLAS; (2) A fused operator of ReLU and $\mathbf{s}\odot(\mathbf{x}\mathbf{W}_{1}^{T})$ with output-side sparsity; (3) A sparse matrix-vector multiplication operator $\mathbf{x}_{1}\mathbf{W}_{2}^{T}$ with input-side sparsity. We adopt the single-step speedup ratios of steps (2) and (3) with these two operators respectively to reflect the practical accurate acceleration potential of sparse LLMs. Refer to Appendix D for implementation details.
4 Experiments
4.1 Experimental Settings
Our training data consists of both language modeling datasets and instruction tuning datasets. For evaluation, we adopt comprehensive benchmarks covering code generation, commonsense reasoning, reading comprehension, and 4 other popular tasks. Refer to Appendix E for more details.
4.2 Overall Results
We apply ProSparse to Swish-activated LLaMA2-7B, LLaMA2-13B Touvron et al. (2023b), and MiniCPM-1B Hu et al. (2024). The obtained sparsely activated models are then compared with their original Swish-activated versions. For comprehensiveness, we also consider ReluLLaMA [^3] https://huggingface.co/SparseLLM/ReluLLaMA-7B, the only open-source ReLU-based LLMs fine-tuned from LLaMA2. All the average sparsity values are computed on the same mixed dataset sampled from training datasets. For more hyper-parameters, see Appendix L and O.
Results are shown in Table 1 (see Appendix G for performance on each independent benchmark). We can draw three conclusions:
(2) Scale Generalizability: The effectiveness of ProSparse consistently holds under three model scales. The promising results on the end-size model (i.e., MiniCPM-1B) reveal the potential of ProSparse as well as activation sparsity on end-user devices, where the inference efficiency of LLMs is significantly emphasized.
(3) Effect of Activation Threshold Shifting: Based on the results without activation threshold shifting (i.e., those with the “” marker), we can demonstrate the effectiveness of this technique in improving the sparsity without compromising performance. Notably, the threshold $t$ must be carefully chosen, see Appendix O.
4.3 Acceleration Effect of Sparsity
Approximate Acceleration Accurate Acceleration
Dense-7B - - - 3.67 1.00 90.55 1.00 82.92 1.00
ReluLLaMA-7B 66.98 90.89 58.95 11.37 3.10 67.12 1.35 63.00 1.32
ProSparse-7B 88.11 93.46 75.24 16.30 4.44 46.66 1.94 55.56 1.49
ProSparse-7B 89.32 92.34 78.75 - - 45.38 2.00 55.05 1.51
Dense-13B - - - 1.92 1.00 131.36 1.00 113.68 1.00
ReluLLaMA-13B 71.56 86.41 71.93 6.59 3.43 69.92 1.88 75.47 1.51
ProSparse-13B 87.97 91.02 77.93 8.67 4.52 55.29 2.38 67.50 1.68
ProSparse-13B 88.80 91.11 78.28 - - 53.78 2.44 66.73 1.70
Approximate Acceleration Algorithm
In this section, we train the activation predictors for each sparse LLM and compute the recalls, predicted sparsity, and actual inference speeds on PowerInfer Song et al. (2023). As the FFN in each Transformer layer has different activation distributions as well as different predictors, the former two metrics are averaged from the results of all layers. Note that MiniCPM-1B has not been tested since PowerInfer does not support its architecture at present. Refer to Appendix F for training details of predictors.
Accurate Acceleration Algorithm
4.4 Analysis and Discussion
Q1: What is the effect of $L_{1}$ regularization and its trend of increase?
For this question, we consider two regularization-free ReLUfication baselines: vanilla ReLU Zhang et al. (2024) and shifted ReLU Mirzadeh et al. (2023). Both only include the substitution of activation functions with ReLU (i.e., $\max(x,0)$) and shifted ReLU (i.e., $\max(x-b,0)$, where $b$ is a tunable bias) respectively.
First, we consider the training dynamics of the above two baselines and ProSparse, as shown in Figure 2. The setting “Fixed $L_{1}$” is a reference setting with a constant regularization factor. Clearly, the training stages with increasing sparsity only include those with regularization applied, namely the whole “Fixed $L_{1}$”, the warmup stage, and the incremental stages of ProSparse. Therefore, among the settings involved, the trend of sparsity is incremental only if non-zero $L_{1}$ regularization is applied [^5] We did not reproduce the flat sparsity trend claimed in the paper of Shifted ReLU Mirzadeh et al. (2023).. Neither vanilla ReLU nor shifted ReLU can push for higher sparsity without regulatization.
However, concerns may naturally arise about the performance, as the additional $L_{1}$ loss term can unavoidably influence the optimization of the language modeling target. For this problem, we evaluate the above methods given different numbers of training tokens. Through experiments (see Appendix H), while a performance gap exists between ProSparse and two baselines given limited 34.60B tokens, it obtains comparable performance when sufficient 89.13B tokens are provided and thus the regularization can increase more smoothly, with a final sparsity value close to the limited-token setting. Therefore, $L_{1}$ regularization can reach far higher activation sparsity and maintain comparable performance to regularization-free methods with a sufficiently smooth increase trend of the factor, at the cost of an acceptable rise in training tokens (i.e., 54.53B, only 2.73% of the 2T tokens used to pre-train LLaMA2 Touvron et al. (2023b)).
Q2: How to reach a target activation sparsity value?
Considering the additional training costs needed to reach higher sparsity, a common requirement lies in how to manipulate ProSparse to reach a desired sparsity given a limited computation budget. The key challenge is the search for suitable regularization factors. To this end, we manage to find the quantitative relationship between the final activation sparsity and the regularization factors to avoid the empirical hyper-parameter search.
Specifically, we select checkpoints at different training stages of ProSparse-7B (see Table 6), apply a constant regularization factor $\lambda_{S}$, and then resume training for sufficient steps (i.e., no less than 4,000 steps) as the last regularization stage. With the same accumulated training token numbers as ProSparse-7B, we can obtain different activation sparsity by tuning the value of $\lambda_{S}$. The results shown in Figure 3 provide two observations: (1) The final activation sparsity is mainly dependent on the last-stage regularization factor $\lambda_{S}$ when $\lambda_{S}$ is relatively large (e.g., $\lambda_{S}>=10^{-2}$ for ProSparse-7B). (2) The activation sparsity shows a negative exponential relationship with $\lambda_{S}^{\alpha}$. For ProSparse-7B, specifically, the sparsity approximates $100-\exp(-1.76\cdot\lambda_{S}^{0.30}+3.49)$ (i.e., the red fitted curve in Figure 3). In summary, to reach a given relatively high sparsity level (e.g., sparsity larger than 80%, satisfying $\lambda_{S}>=10^{-2}$), the only thing needed is to control the regularization factor $\lambda_{S}$ of the final progressive regularization stage. Therefore, given the fixed model size, ProSparse is a highly controllable ReLUfication method in terms of sparsity adjustment.
Q3: Is progressive sparsity regularization effective?
If the activation sparsity mainly depends on the final-stage regularization factor, why should we increase the factor progressively? The answer lies in the performance concern. To substantiate the effectiveness of progressive sparsity regularization, the second step of ProSparse, we conduct ablation studies by making the regularization factor constant throughout the training process after activation function substitution. By setting the factor to $0.1$, we obtain a model with activation sparsity of 88.62%, slightly lower than ProSparse-7B (89.32%). However, with the same number of training tokens, this model only has an average performance of 36.34%, considerably lower than ProSparse-7B (38.46%). Similarly, for the 13B setting, we obtain a model with the comparable sparsity of 88.96% and an average performance of 42.85%, lower than ProSparse-13B (see Appendix I). Therefore, progressive sparsity regularization is indispensable in mitigating the performance degradation caused by ReLUfication.
Q4: How to SFT sparsely activated models?
It is non-trivial to apply SFT to sparsely activated models obtained by ProSparse. Our practice of training ProSparse-1B can provide some experience: SFT can be applied to sparsely activated models obtained by ProSparse with a well-chosen regularization factor, and this factor for SFT is empirically smaller than $\lambda_{S}$ to accommodate newly injected knowledge and avoid performance degradation. See Appendix J for more details and observations.
Q5: How does the sparsity distribute?
Another interesting observation is the imbalanced sparsity distributions among distinct datasets and layers. Specifically, the activation sparsity of ProSparse models is higher on more formatted instruction tuning datasets and higher layers (i.e., layers closer to outputs). More detailed analyses are provided in Appendix M and N.
5 Conclusion
In this work, we propose ProSparse, an effective ReLUfication method for introducing and enhancing intrinsic activation sparsity from non-ReLU LLMs with comparable performance. Extensive experiments demonstrate the effectiveness of ProSparse and its practical values in inference acceleration with various algorithms. Deeper analyses concerned with certain ProSparse techniques, model properties, and SFT issues further substantiate the practicality of ProSparse and provide valuable insights.
Broader Impacts
This paper presents a simple and effective method, ProSparse, for introducing and enhancing ReLU-based intrinsic activation sparsity into non-ReLU LLMs. There may exist many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Limitations
Firstly, more comprehensive studies on huge-scale models (e.g., 70B or more) should be included in the future given sufficient computing resources. Moreover, we only focus on the sparsity-based acceleration of step (2) and step (3) of FFN, leaving a considerable ratio of LLM computation unoptimized. Actually, there already exist preliminary works in the sparsification of the attention layers Shen et al. (2023); Wortsman et al. (2023). Methods such as pruning and low-rank decomposition may also be helpful in optimizing the FFN step (1) Ji et al. (2024). For future works, we will continue to explore how to introduce and enhance sparsity in the attention layer as well as the acceleration issue of the FFN step (1).
Acknowledgments
This work is supported by National Natural Science Foundation of China (No. 62236004, No. 62236011, No. 62302479), China Postdoctoral Science Foundation (2023M733566), and Institute Guo Qiang at Tsinghua University.
Appendix A Extended Related Works
$L_{1}$ Regularization
In statistical learning such as linear regression, $L_{1}$ regularization has been long adopted as a classical and effective technique for sparsification Tibshirani (1996); Hastie et al. (2009). With the advent of deep learning, researchers also explore paradigms of applying $L_{1}$ regularization to neural networks. One prominent usage is model pruning Cheng et al. (2017). Specifically, a term of loss calculated as the $L_{1}$ norm of the sparsification target is added to the optimization target function to prompt sparse weights for faster computation. This has helped acceleration in various conventional neural networks Han et al. (2015); Zhao et al. (2016); Wen et al. (2016); Scardapane et al. (2017); Ma et al. (2019); Wang et al. (2019) as well as Transformer-based models Zhu et al. (2021); Prasetyo et al. (2023). Inspired by these works, some researchers also try to adopt $L_{1}$ regularization for activation sparsity, mainly in ReLU-activated convolutional networks Georgiadis (2019); Kurtz et al. (2020) and Transformer-based architectures Li et al. (2022).
To the best of our knowledge, ProSparse is the first work using a dynamic $L_{1}$ regularization factor for prompting activation sparsity in neural networks. By contrast, a majority of the former works adopt fixed factors. For more adaptive control, some of them introduce group regularization Yuan and Lin (2006), namely using different factors for different parameter groups. Nevertheless, without dynamic factors, these paradigms can cause a substantial shift in activation distribution and thus potentially risk performance degradation. The work most related to ProSparse is IncReg Wang et al. (2019), which introduces incremental regularization factors that change for different parameter groups at each iteration. While they focus on the pruning of convolutional networks, ProSparse handles a distinct scenario of prompting activation sparsity in Transformer-based LLMs and adopts a completely different strategy consisting of a progressively incremental factor.
Difference between Activation Sparsity and Pruning
Generally, pruning realizes sparsity by removing certain elements (e.g., neurons, weights, or structured blocks) in LLMs. However, the sparsity introduced by pruning is statically limited to model weights and independent of the inputs. High static sparsity is often accompanied by considerable performance degradation compared to the original dense model Frantar and Alistarh (2023); Xia et al. (2023). By contrast, activation sparsity is dynamically determined by the input data and thus potentially compromises less model capacity and downstream task performance.
Appendix B Detailed Algorithm for Progressive Sparsity Regularization
@@
The detailed algorithm for scheduling the factor $\lambda$ in progressive sparsity regularization is listed in Algorithm 1.
Algorithm 1 Progressive factor scheduling adopted in ProSparse
@@
Appendix C Extented Introduction of Approximate Acceleration Algorithms
Existing approximate algorithms are mostly dependent on activation predictors, which are small neural networks to predict the intermediate activations $\mathbf{x}_{1}$ based on the input hidden states $\mathbf{x}$ Liu et al. (2023); Song et al. (2023). If one element at a specific position of $\mathbf{x}_{1}$ is predicted to be zero, then all the computations associated with this position can be saved with little or no hardware resources allocated. This is the key mechanism with which approximate algorithms realize acceleration.
Nevertheless, such a predictor-dependent acceleration effect is largely dependent on the performance of the pre-trained activation predictors. For example, a typical bad case is that an actually activated element in $\mathbf{x}_{1}$ is predicted to be inactivated. This can bring about negative results including unwise hardware resource allocation and erroneously ignored intermediate logits, which limits the practical speedup ratios and even causes inference inaccuracies. Therefore, a sparse LLM can gain more benefits from approximate algorithms if its activation distribution is more predictable.
To test a sparse LLM’s practical acceleration value with approximate algorithms, we involve the predictability of its activation distribution, which is evaluated by the performance of its specifically pre-trained activation predictor. This involves two key metrics: the activation recall and the predicted sparsity.